# Can LLMs Recognize Toxicity? Definition-Based Toxicity Metric

### v3

@@ 6.2 LLM Evaluator
Evaluating the generated output of an NLP model can be broadly divided into two categories: lexical-based metrics and semantic-based metrics. The first category encompasses metrics that rely on lexical features of references such as BLEU (Papineni et al., 2002), Rouge (Lin, 2004), chrF++ (Popović, 2017), and spBLEU (Goyal et al., 2022). The second category involves metrics that consider semantic aspects such as BertScore (Zhang et al., 2020), COMET (Rei et al., 2020), Clipscore (Hessel et al., 2022). Our LATTE’s score can be interpreted as a discrete score of the decoder akin to encoder model scores as BertScore (Yuan et al., 2021). Recently, numerous studies have highlighted the possibility of LLMs functioning as evaluators in different domains, namely machine translation quality (Kocmi and Federmann, 2023), summarization quality (Liu et al., 2023), dialogue quality (Plátek et al., 2023; Lin and Chen, 2023), and reasoning quality (Yao et al., 2023).
@@ 7 Conclusion
Despite the rapid advancements in the AI field, metrics related to toxicity remain in a state of stagnation. While recent research evaluate semantic areas such as naturalness by using LLMs, such methodologies should not be applied without caution in the field of toxicity. In our research, we propose a LLM-based evaluation metric that considers diverse LLM factors. We find that current LLMs are reliable within confined factors. Our methodology demonstrates superior performance to existing bertology approaches in verified factors. We show that LLMs adeptly follow the provided definition instructions, and our evaluation prompt is robust against diverse perturbation and applicable to various LLMs. Nevertheless, LLMs struggle to handle argumentative factors, indicating that further research is needed in such area.
@@ 8 Limitations
Our methodology requires computational costs due to the substantial size of LLMs compared to traditional models. Nevertheless, its performance and flexibility are efficient enough to be deployed and could be naturally mitigated with the advent of lightweight LLMs in the future. During the process of constructing benchmarks, there may be biases in our selection. Therefore, we have endeavored to utilize a diverse range of existing datasets and have made efforts to ensure a fair comparison to the best of our knowledge. Due to both the limited mathematical capabilities of LLMs and the lack of existing benchmarks capable of measuring the extent of toxicity on a continuous scale, we confined our experimental setup to a binary framework. We anticipate that these limitations will be naturally resolved as the fundamental abilities of LLMs improve and more datasets are released.
9 Discussion and Future Work
Our experimental results reveal that performance varies based on the content and format of the evaluation prompt. Recent research (Bertolazzi et al., 2023), demonstrates that LLMs tend to behave more like children than adults through a game of 20 questions. It implies that we have to provide detailed information regarding toxicity and relevant contexts for controlling LLMs. To provide such a detailed information during the period of LLMs’ proliferation, the academic community needs to engage in discussions to set up the minimum standard for toxicity, not just keyword-based and trigger-prompts approach, in order to steer LLMs to be prosocial. Furthermore, as simple prompts for measuring toxicity is comparable to those of existing encoder-based scores, this results pose the questions whether LLMs may possess certain inherent perspectives on toxicity or just pretend to possess. Derived from such questions, we need to discuss about whether they can self-improve based on such cognizance.
9.1 Measuring Ethical Perspectives
We further define a new factor of toxicity, referred to as justice, derived from the seven AI trustworthiness factors Jobin et al. (2019). The concept of justice is technically interpreted in Hendrycks et al. (2021). From an ethical perspective, individual values are segmented into virtue, deontology, and utilitarianism categories. Utterances that are not explicitly demeaning or partial but conflict with an individual’s ethical values may still cause discomfort. We therefore categorize such instances as ’weak toxic.’ We collectively nominate these factors as Ethical Preference. However, reaching a consensus on whether utterances are toxic is more complicated than simply identifying demeaning or partial content.
In Appendix B.3, we evaluate the ethical preference capabilities of LLMs. Similar to our sanity check process, we assess their ability to discern ethical perspectives, yet LLMs fail to align with ethical perspectives. To inform LLMs with relevant theories, we refer to existing research (Zhou et al., 2023). however, the empirical results show that their ethical capabilities are unreliable.
Nevertheless, it remains unclear whether LLMs truly lack an understanding of ethical principles, as there are very few benchmarks available to assess whether these models possess an awareness of ethical perspectives. Interpreting situations from diverse ethical perspectives often fails to yield a consensus, owing to the inherently subjective nature of ethics. Such subjectivity complicates the creation of objective benchmarks for testing these capabilities.
Moreover, we empirically observe that although LLMs fail to score generated outputs, they can construct reasonable responses when instructed to generate replies based on certain ethical perspectives. Recent studies have assessed the ethical capacity of models by presenting them with moral dilemmas to evaluate their ethical awareness. However, as indicated by our prompt ablation study, LLMs are influenced by multiple factors, and mere instructions may not ensure ethical awareness. This highlights the necessity for further research into developing effective methods to measure ethical awareness in LLMs.
@@ 10 Ethical Statements
Defining what constitutes toxicity and interpreting it is intricate. Therefore, we try to establish relevant concepts based on previous articles, and set up benchmarks to measure responses by incorporating previous toxicity research as much as possible. We reflect various AI ethical guidelines, philosophical domains, and engineering domains in a comprehensive manner to define the scope of toxicity in our work concerning toxicity in LLMs. Our research suggests that there is a substantial margin for enhancement and development of moral value assessment, beyond offensiveness and demographic bias. Adopting our metric for value assessment without any sanity-check process causes unintended problems.
@@ B.2 Partiality
In Figure 13 and 14, We obtain negative stereotypes and disambiguous context from BBQ. We test the demographic bias for GPT and Llama by BBQ dataset. Naive means simple QA format prompt, COT for adding think step-by-step, COT2 for adding take a deep breath and think step-by-step, Safe for default llama’s system instruction, and Safe COT for adding COT prompt to Safe prompt. They are all well aware of negative demographic bias.
@@ D.1 Evaluation Prompt’s Factors
Model Type and Content The choice of model is also crucial when deploying LATTE. In the case of ParaDetox evaluation, Llama2-Code shows remarkable results on the evaluation of safe utterance, while Llama2-NLP achieves outstanding performance on the evaluation of toxic. However, it can be seen that Llama2 exhibits higher variance compared to GPT-4 in response to alterations of prompt and dataset. Figure 16 and 17 show the fine-grained results of content for each model. In the Figure 5, Toxic is for toxic utterances, Safe for safe utterances, NLP for natural instructions, and Code for code format instructions. Naive is for simple instruction, Ref for antonym, COT for reasoning content, Sans and Kor for multilingual content, and Meaning for definition. Details are in Appendix D.
@@
Format Figure 16 and 17 show that the format of the prompt used in LATTE considerably affects performance. The code format generally picks out safe utterances with better precision than the NLP format with Llama2 70B, and the NLP format results in better accuracy in classifying toxic utterances than the code format with Llama2 70B.
@@ F.1 Paradetox & Prosocial Dialogue
In Table 5, the primary elements to focus on are the ‘Toxic’ accuracy for comparison as the primary focus of existing studies is the ’Toxic’ row of Table 13. Indeed, the results using Llama2 70B and GPT-4 in the context of Toxic significantly surpass PerspectiveAPI (57.8[PerspectiveAPI], 81.8/72.2[LATTE]). Notably, it is observed that PerspectiveAPI Proso. Toxic fails to detect toxicity, suggesting a potential inability to discern contextual toxic information, even in a binary setting. Consequently, sequences that are actually toxic are often blindly misclassified as safe. This discrepancy is not adequately captured by ACC, thus we introduced the F1 score for additional demonstration. According to evaluations based on F1 scores, LATTE significantly outperforms PerspectiveAPI, indicating that PerspectiveAPI faces challenges in detecting discriminatory contexts or subtle forms of aggression in routine conversations.
@@ F.5 Robustness of LATTE under perturbations
We perturb the LATTE prompt as shown in Table 20 on the Prosocial Dialog dataset and the Paradetox dataset. Results show that modifying the temperature causes insignificant variations to LATTE’s performance. In addition, perturbing the prompt has a more significant impact on performance within the Prosocial Dialog dataset, compared to the Paradetox dataset. Texts labelled as toxic in the ParaDetox dataset contain profanities in general. In contrast, texts labelled in Prosocial may not contain profanities, but may implicitly go against prosocial norms, thus making the evaluation task more complex. Those difference contributes to the difference of perturbing effects. Still, the deviations from the original performance caused by perturbations does not exceed 4% when the definition of toxicity is not ambiguous. Modifying the separator between contents also does not cause significant changes to the performance, as illustrated in Sclar et al. (2023). We can thus see that prompts with detailed descriptions can lead to better robustness.

### v4

@@ 6.2 LLM Evaluator
Evaluating the generated output of an NLP model can be broadly divided into two categories: lexical-based metrics and semantic-based metrics. The first category encompasses metrics that rely on lexical features of references such as BLEU (Papineni et al., 2002), Rouge (Lin, 2004), chrF++ (Popović, 2017), and spBLEU (Goyal et al., 2022). The second category involves metrics that consider semantic aspects such as BertScore (Zhang et al., 2020), COMET (Rei et al., 2020), UMIC (Lee et al., 2021), Clipscore (Hessel et al., 2022). Our LATTE’s score can be interpreted as a discrete score of the decoder akin to encoder model scores as BertScore and BartScore (Yuan et al., 2021). Recently, beyond such encoder-based models, numerous studies have highlighted the possibility of LLMs functioning as evaluators in different domains, namely machine translation quality (Kocmi and Federmann, 2023; Lee et al., 2024), summarization quality (Liu et al., 2023), dialogue quality (Plátek et al., 2023; Lin and Chen, 2023; Hwang et al., 2024), and reasoning quality (Yao et al., 2023).
@@ 7 Conclusion
Despite the rapid advancements in the AI field, metrics related to toxicity remain in a state of stagnation. Recently, there has been an increase in research that evaluates semantic areas such as naturalness by using LLMs. However, such methodologies should not be applied without caution in the field of toxicity. In our research, we propose a toxicity investigation framework and an evaluation metric that considers diverse LLM factors. We find that current LLMs are reliable within confined factors. Recent research (Bertolazzi et al., 2023), demonstrates that LLMs tend to behave more like children than adults. It implies that we have to provide detailed information regarding toxicity and relevant contexts for controlling LLMs. To provide such a detailed information during the period of LLMs’ proliferation, we need to discuss the minimum standard for toxicity, not just keyword-based and trigger-prompts approach, in order to steer LLMs to be prosocial.
@@ 8 Limitations
Our methodology requires computational costs due to the substantial size of LLMs compared to traditional models. Nevertheless, its performance and flexibility are efficient enough to be deployed and could be naturally mitigated with the advent of lightweight LLMs in the future. During the process of constructing benchmarks, there may be biases in our selection. Nevertheless, we have endeavored to utilize a diverse range of existing datasets and have made efforts to ensure a fair comparison to the best of our knowledge. Due to both the limited mathematical capabilities of LLMs and the lack of existing benchmarks capable of measuring the extent of toxicity on a continuous scale, we confined our experimental setup to a binary framework. We anticipate that these limitations will be naturally resolved as the fundamental abilities of LLMs improve and more datasets are released. Lastly, our methodology struggles to conduct value assessments related to addressing moral issues and preference alignment beyond normative contents. Continuous research is necessary to progress in this direction.
@@ 9 Ethical Statements
Defining what constitutes toxicity and interpreting it is intricate. Therefore, we try to establish relevant concepts based on previous articles, and set up benchmarks to measure responses by incorporating previous toxicity research as much as possible. We reflect various AI ethical guidelines, philosophical domains, and engineering domains in a comprehensive manner to define the scope of toxicity in our work concerning toxicity in LLMs. Our research suggests that there is a substantial margin for enhancement and development of moral value assessment, beyond offensiveness and demographic bias. Adopting our metric for value assessment causes unintended problems.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics]. This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)]. K. Jung is with ASRI, Seoul National University, Korea. The Institute of Engineering Research at Seoul National University provided research facilities for this work.
@@ B.2 Partiality
In Figure 11 and 12, We obtain negative stereotypes and disambiguous context from BBQ. We test the demographic bias for GPT and Llama by BBQ dataset. Naive means simple QA format prompt, COT for adding think step-by-step, COT2 for adding take a deep breath and think step-by-step, Safe for default llama’s system instruction, and Safe COT for adding COT prompt to Safe prompt. They are all well aware of negative demographic bias.
| Model | Llama2 | GPT | ||||
| Size | 7B | 13B | 70B | 3 | 3.5 | 4 |
| Acc | 0.785 | 0.777 | 0.796 | 0.855 | 0.950 | 0.957 |
In sum, all LLMs are well aware of demographic bias even without few-shot examples according to Table 5.
@@ D.1 Evaluation Prompt’s Factors
Model Type and Content The choice of model is also crucial when deploying LATTE. In the case of ParaDetox evaluation, Llama2-Code shows remarkable results on the evaluation of safe utterance, while Llama2-NLP achieves outstanding performance on the evaluation of toxic. However, it can be seen that Llama2 exhibits higher variance compared to GPT-4 in response to alterations of prompt and dataset. Figure 15 and 16 show the fine-grained results of content for each model. In the Figure 4, Toxic is for toxic utterances, Safe for safe utterances, NLP for natural instructions, and Code for code format instructions. Naive is for simple instruction, Ref for antonym, COT for reasoning content, Sans and Kor for multilingual content, and Meaning for definition.
Specifically, in Figure 18, the provision of a Meaning of toxicity (def) significantly enhances performance in toxicity detection compared to the omission of toxicity definition (no_def). This improvement is consistent across all models and evaluation datasets, with the exception of GPT4-u-proso, which shows unchanged performance. The Prosocial Dialogue is particularly more complex than Paradetox, as it challenges the consensus on definitive toxicity. Thus, when the LATTE metric is applied with a specific definition, it effectively captures and measures toxicity.
Guideline: “It’s wrong to take part in violence.”
Context: “He made up a conspiracy theory that was far from the truth.”
Guideline: ‘It’s okay to be upset with someone for being rude to you.”
@@
Format Figure 15 and 16 show that the format of the prompt used in LATTE considerably affects performance. The code format generally picks out safe utterances with better precision than the NLP format with Llama2 70B, and the NLP format results in better accuracy in classifying toxic utterances than the code format with Llama2 70B.
@@ F.1 Paradetox & Prosocial Dialogue
We use safe as an auxiliary metric to check whether the model blindly considers sequences toxic, and the results related to safe are in Table 12. The primary elements to focus on are the ‘Toxic’ accuracy for comparison as the primary focus of existing studies is the ’Toxic’ row of Table 12. Indeed, the results using Llama2 70B and GPT-4 in the context of Toxic significantly surpass PerspectiveAPI (57.8[PerspectiveAPI], 81.8/72.2[LATTE]). Notably, it is observed that PerspectiveAPI Proso. Toxic fails to detect toxicity, suggesting a potential inability to discern contextual toxic information, even in a binary setting. Consequently, sequences that are actually toxic are often blindly misclassified as safe. This discrepancy is not adequately captured by ACC, thus we introduced the F1 score for additional demonstration. According to evaluations based on F1 scores, LATTE significantly outperforms PerspectiveAPI, indicating that PerspectiveAPI faces challenges in detecting discriminatory contexts or subtle forms of aggression in routine conversations.
@@ F.5 Robustness of LATTE under perturbations
| Type | Modification | Prosocial | Paradetox | |||
| Toxic | Safe | Toxic | Safe | |||
| 0 | Format | Casing | -2.4 | +0.8 | +1.2 | -0.8 |
| Spacing | +2.8 | -1.2 | +1.2 | -0.8 | ||
| Seperator | +2.0 | -1.2 | +2.4 | -1.2 | ||
| Period Delete | +0.0 | +0.8 | +2.8 | +0.0 | ||
| Definition | Paraphrase | -7.6 | +3.6 | +2.4 | -0.8 | |
| + Period Delete | -6.0 | +2.4 | +3.2 | -1.2 | ||
| 0.5 | Format | Casing | -1.6 | +1.2 | -2.4 | -0.8 |
| Spacing | +1.6 | -2.0 | +0.4 | -0.8 | ||
| Seperator | +2.0 | -2.8 | +0.8 | -1.6 | ||
| Period Delete | -0.8 | +1.2 | +2.0 | -1.2 | ||
| Definition | Paraphrase | -8.8 | +3.2 | +3.2 | -2.0 | |
| + Period Delete | -5.6 | +3.2 | +4.0 | -2.4 | ||
| 1.0 | Format | Casing | -3.6 | +0.8 | -2.4 | +0.4 |
| Spacing | +1.6 | +0.0 | +1.2 | -0.8 | ||
| Seperator | +2.4 | -0.4 | -0.4 | -0.8 | ||
| Period Delete | +0.4 | +1.6 | +0.4 | -1.6 | ||
| Definition | Paraphrase | -8.4 | +4.8 | +1.6 | -1.2 | |
| + Period Delete | -8.8 | +2.4 | +2.2 | -1.2 | ||
Our experimental results show that introducing perturbations to definition prompts does not make significant variance, as shown in Table 17 Format rows. Sufficiently providing few-shot prompts to resolve the variance problem, our LATTE becomes more robust in Table 19. These results demonstrate that creating an evaluation dataset is still essential for accurately measuring customized toxicity that is not influenced by prompt perturbations.
@@
In detail, we perturb the LATTE prompt as shown in Table 20 on the Prosocial Dialog dataset and the Paradetox dataset. Results show that modifying the temperature causes insignificant variations to LATTE’s performance. In addition, perturbing the prompt has a more significant impact on performance within the Prosocial Dialog dataset, compared to the Paradetox dataset. Texts labelled as toxic in the ParaDetox dataset contain profanities in general. In contrast, texts labelled in Prosocial may not contain profanities, but may implicitly go against prosocial norms, thus making the evaluation task more complex. Those difference contributes to the difference of perturbing effects. Still, the deviations from the original performance caused by perturbations does not exceed 4% when the definition of toxicity is not ambiguous. Modifying the separator between contents also does not cause significant changes to the performance, as illustrated in Sclar et al. (2023). We can thus see that prompts with detailed descriptions can lead to better robustness.